OpenStudy (anonymous):

@ryanwhite499 Please help

OpenStudy (anonymous):

http://prntscr.com/8edkqe

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. −4x+5(x−2)=−16+10 Step 1: Simplify both sides of the equation. −4x+5(x−2)=−16+10 Simplify: (Show steps) x−10=−6 Step 2: Add 10 to both sides. x−10+10=−6+10 x=4 Answer: x=4